Question 3 1 pts Two waves with the same frequency and wavelength but with different amplitudes are added. If A1 = 2A2 and th.
Answer #1

solution (Question 3) resultant amplitude - A then A = VARTA 2 +24, Azcose =A7+ A3 +2.442.603 1800 = J A²+A²-2002 = VAI A272